OpenStudy (anonymous):

6561^(2n)=729^(n+2) ummm N?

OpenStudy (anonymous):

U know logarithm?

OpenStudy (anonymous):

um no..

OpenStudy (anonymous):

\[9^{4{\times}2n}=9^{4{\times}(n+2)}\]\[4\times(2n)=4\times(n+1)\]\[8n=4n+4\]\[4n=4\]\[n=1\]

OpenStudy (anonymous):

so my n would equal 2?

OpenStudy (anonymous):

you are asking me that " Is the sky green?" I just say n=1

OpenStudy (anonymous):

well. that was rude. 1 isnt one of my options

OpenStudy (anonymous):

oh sorry it should be 4*(2n)=3*(n+1) w/c gives n = 3/5

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

not really but im like math awkward

OpenStudy (anonymous):

ok 6561 = 9^4 729 = 9^3 Can u insert them in there respective place?

OpenStudy (anonymous):

not really, tis why im here and tis why i failed this class last year.

OpenStudy (anonymous):

guess you're gonna fail again ahahhhhhhhhhhhhhaahahaha

OpenStudy (anonymous):

6561^(2n)=729^(n+2) log (6561^2n) = log(729^[n+2]) 2nlog(6561) = (n+2)log(729) finish it.

OpenStudy (anonymous):

that was so mean im a little butt hurt right now.

OpenStudy (anonymous):

2nlog(6561) = nlog(729) + 2log(729) 2nlog (6561) - nlog (729) = 2log (729) n(2log[6561] - log [729]) = 2log (729) n = 2log (729) / [2log(6561) - log (729)] There you go honey.

OpenStudy (anonymous):

okay thanks boo.

OpenStudy (anonymous):

remember, i don't like pickles on my big mac.

OpenStudy (anonymous):

oh i thought you said you like extra pickles.. damn

OpenStudy (anonymous):

hahahaha, so after you take a base of 27 from each logarithm, your final answer will be 6/5

OpenStudy (anonymous):

clean up on isle 4 please?

OpenStudy (anonymous):

6/5? oh goodness. im going to fail this is going to suck. thank you

OpenStudy (anonymous):

hold up.

OpenStudy (anonymous):

you spilt something in isle 4

OpenStudy (anonymous):

oh heeeellllll no

OpenStudy (anonymous):

what do you not understand about my steps?

OpenStudy (anonymous):

i simply took the logarithm base 27 from both sides, applied the power rule, factored and isolated for n.

OpenStudy (anonymous):

no, its not you really. its me. i'm math challenged. i question stuff too much

OpenStudy (anonymous):

wait that kinda made sense!

OpenStudy (anonymous):

oh who's a big boy now?